An antiproton has the mass of a proton but a charge of -e. If a proton and an antiproton orbited each other, how far apart would they be in the ground state of such a system? Why might you think such a system could not occur?

1
Expert's answer
2021-12-06T09:39:59-0500

Find the Bohr radius for such a hydrogen-like system:


"a_0=\frac{\epsilon_0h^2}{\pi m_pe^2}=2.88·10^{-14}\text{ m}."

Such a system could not occur because the radius of a proton is 0.84·10-15 m and they would annihilate.